# Quantum circuit complexity and unsupervised machine learning of topological order

### v1

@@ Abstract
Inspired by the close relationship between Kolmogorov complexity and unsupervised machine learning, we explore quantum circuit complexity, an important concept in quantum computation and quantum information science, as a pivot to understand and to build interpretable and efficient unsupervised machine learning for topological order in quantum many-body systems. To span a bridge from conceptual power to practical applicability, we present two theorems that connect Nielsen’s quantum circuit complexity for the quantum path planning between two arbitrary quantum many-body states with fidelity change and entanglement generation, respectively. Leveraging these connections, fidelity-based and entanglement-based similarity measures or kernels, which are more practical for implementation, are formulated. Using the two proposed kernels, numerical experiments targeting the unsupervised clustering of quantum phases of the bond-alternating XXZ spin chain, the ground state of Kitaev’s toric code and random product states, are conducted, demonstrating their superior performance. Relations with classical shadow tomography and shadow kernel learning are also discussed, where the latter can be naturally derived and understood from our approach. Our results establish connections between key concepts and tools of quantum circuit computation, quantum complexity, and machine learning of topological quantum order.
@@ I Introduction
Recent progress has shown the potential of machine learning for classifying topological phases of matter Carleo and Troyer (2017); Carrasquilla and Melko (2017); Van Nieuwenburg et al. (2017); Zhang and Kim (2017); Liu and van Nieuwenburg (2018); Zhang et al. (2018); Sun et al. (2018); Carleo et al. (2019); Ming et al. (2019); Rem et al. (2019); Lian et al. (2019); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020); Long et al. (2020); Greplova et al. (2020); Balabanov and Granath (2020); Yu and Deng (2021); Mendes-Santos et al. (2021); Leykam and Angelakis (2023); Smolina et al. (2024); Das Sarma et al. (2019). In particular, the supervised approach to quantum many-body physics problems, which can resort to well established theories of learnability and generalizability Valiant (1984); Mohri et al. (2018); Jacot et al. (2018), has resulted in the formulation of rigorous guarantees Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). However, in many practical scenarios, where a priori knowledge of the different phases of matter (and accordingly their labels) does not exist, only unsupervised machine learning can provide access to the desired information. Despite the practical relevance of unsupervised learning, a rigorous theory, tailored to topological quantum order and offering both interpretability and generalizability, has not yet been fully explored.
@@
One prevalent approach to unsupervised learning is based on kernel methods Mohri et al. (2018); Jacot et al. (2018); Rodriguez-Nieva and Scheurer (2019), where feature vectors of the data set are encoded into a kernel in terms of their overlaps or similarities, followed by a nonlinear dimensionality reduction for clustering in a low-dimensional representation space, e.g., the kernel principal component analysis (PCA) Schölkopf et al. (1998); Mika et al. (1999) or the diffusion map algorithm Coifman et al. (2005); Nadler et al. (2006). In the context of topological quantum order, the task then is to formulate a kernel that encodes the nonlinear features of topological quantum states and represents topological similarities or distances by exploiting the primitive notion of topological equivalence. Topologically equivalent quantum states, by definition, can be continuously transformed into each other via a smooth unitary transformation generated by a local Hermitian operator with bounded operator norm Zeng et al. (2015). In this spirit, a kernel based on a path-finding algorithm has been proposed in Scheurer and Slager (2020) and shown good performance in clustering symmetry-protected band topologies, which are characterized by short-range quantum entanglement. Other topological similarity measures, that focus on the closing of the spectral gap at topological transitions, have also been proposed and demonstrated Che et al. (2020); Long and Zhang (2023).
@@
The inherent problems of previous kernel formulations call for a new kernel ansatz, an ansatz which is able to capture the topological distance and entanglement patterns at different scales while at the same time remaining interpretable and efficiently computable (see Fig. 1 for a schematic of the current state regarding the application of machine learning to the classification of topological phases of matter). To this end, we here adopt the alternative but equivalent perspective that two topologically equivalent multi-qubit states can be transformed into each other via a constant-depth quantum circuit composed of geometrically local quantum gates Zeng et al. (2015); Zeng and Wen (2015), i.e., the smooth unitary transformation that connects them can be decomposed into a shallow quantum circuit. This implies that it is algorithmically trivial or cheap to generate one quantum state given the access to another in the same topological class by fully exploiting the shared structure and entanglement patterns between them.
@@
Figure 1: Landscape of machine learning topological phases of matter. Supervised learning requires prior information or labels, whose efficiency guarantees have been extensively investigated Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). Under the constraint of geometric locality Lewis et al. (2024); Rouzé et al. (2024) or with bounded number (No.) of parameters (param.) Che et al. (2024), provably efficient machine learning of quantum many-body systems has been established, e.g., in the framework of the probably approximately correct (PAC)-learnable Valiant (1984); Che et al. (2024). Supervised learning is conceptually more straightforward due to the existence of well-established statistical learning theories. In contrast, unsupervised learning, whose theory is more complicated and less addressed, becomes practically relevant in case of no access to prior labels. The symmetry-protected topological (SPT) order features short-range entanglement, while the topological quantum order exhibits long-range entanglement. The unsupervised learning with the path-finding algorithm Scheurer and Slager (2020) and kernels focusing on closing gaps at topological transitions Che et al. (2020); Long and Zhang (2023), respectively, have shown good performance and interpretability for SPT orders or band topologies. However, the direct generalization of these methods to more complex quantum systems with increased complexity and quantum entanglement faces challenges, due to the intrinsic hardness of identifying (or even the undecidability Cubitt et al. (2015) of) the spectral gap for generic quantum many-body Hamiltonians (See more elaborations and references in the introduction of the main text). Note that the neural network representation of quantum states and reinforcement learning for quantum systems are not included in this landscape.
@@
The correspondence between topological equivalence of quantum states and shallow quantum circuit transformations indicates that, a complexity measure may be related to the topological distance. This is similar to the situation with Kolmogorov complexity, which represents an informational distance between two strings or objects Bennett et al. (2010), and may offer a basis for theoretically optimal solution to unsupervised machine learning Ily (See more details in the following section). We then assert that, the (conditional) generation and prediction of quantum states with the minimal quantum circuit cost may offer a theoretically optimal solution to the unsupervised learning of topological order. The complexity of the generating circuit can be interpreted as a quantum informational distance measure that captures the entanglement change at different scales. Specifically, we propose to use the quantum circuit complexity (QCC) of the unitary path connecting two quantum states as a a suitable distance measure that can be used to construct a kernel for the unsupervised learning of topological order. While the QCC is already a central concept in quantum information theory Chi-Chih Yao (1993); Brandão et al. (2021); Haferkamp et al. (2022); Yi et al. (2024); Liu et al. (2020a), quantum field theory Chapman et al. (2018); Hackl and Myers (2018); Khan et al. (2018) and quantum gravity Susskind (2014); Stanford and Susskind (2014); Brown et al. (2016a, b, 2023), its connection with, and relevance for, the unsupervised machine learning of topological phases of matter has remained mostly unexplored.
@@
In this work, we use the QCC as a pivotal tool to understand and construct unsupervised machine learning of topological order for quantum many-body systems, in an interpretable and efficient manner. Although the exact QCC is generically intractable, we identify various quantities which are upper bounded by the QCC while remaining implementable, and which thus can be efficiciently utilized for classifying topological quantum many-body phases. Specifically, we establish entanglement-based and fidelity-based kernels, respectively. Numerical experiments with a bond-alternating XXZ qubit chain and Kitaev’s toric code model, respectively, demonstrate the excellent performance of these kernels.
II Quantum path planning, topological order and quantum circuit complexity
@@
As mentioned above, the unsupervised machine learning of band topology has been extensively investigated. Band topologies, or the more general symmetry-protected topological (SPT) orders, such as those exhibited by topological insulators, are characterized by the absence of long-range quantum entanglement. This implies that, if the constraint of preserving the symmetry is dropped, SPT quantum states can be transformed smoothly into product states. In contrast, topological quantum order refers to quantum states with long-range quantum entanglement Zeng et al. (2015); for instance, the ground state of Kitaev’s toric code exhibits topological quantum order that manifests in non-zero topological entanglement entropy Kitaev and Preskill (2006); Levin and Wen (2006) and global logical operations Kitaev (2003), while SPT phases are confined to short-range entanglement. Thus, the task of identifying topological quantum order with unsupervised learning presupposes to construct kernel (similarity) functions with feature maps that capture quantum entanglement at a variety of scales. For this purpose, we start by defining the quantum path planning between two multi-qubit quantum states, which lays the ground for our search for a distance measure of clustering topological orders.
@@
[Quantum path planning (QPP).] The QPP problem is defined as: Given two arbitrary pure $n$-qubit quantum states $\rho_{0}$ and $\rho_{1}$, find continuous paths $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$ $(s\in[0,1])$ that connect $\rho_{0}$ and $\rho_{1}$, where the generator $G(s)$ of $U(s)$ is a hermitian operator for all $0\leq s\leq 1$, and $\|\partial_{s}U(s)\|_{\infty}=\|G(s)\|_{\infty}<+\infty$. In general, we have that $U(s)\in SU\left(2^{n}\right)$ and $G(s)\in su\left(2^{n}\right)$, with the unitary given by
@@
where $U(s=0)=\mathrm{Id}$ is the identity and $\cal{P}$} denotes the path ordering operator.
@@
(Topological equivalence Zeng and Wen (2015); Zeng et al. (2015).) We say that two gapped ground states $\rho_{0}$ and $\rho_{1}$ of local Hamiltonians $H_{0}$ and $H_{1}$, respectively, are in the same topological phase if there is a local unitary that maps the two states into each other without closing the energy gap. In other words, there is a smooth path of local Hamiltonians $H(s)$ that connects the two parent Hamiltonians with $H(s=0)=H_{0}$ and $H(s=1)=H_{1}$ for $s\in[0,1]$. The Hamiltonian $H(s)$ is gapped above its ground state $\rho(s)$ for all $s\in[0,1]$, where $\rho(s)$ smoothly connects $\rho_{0}$ and $\rho_{1}$, which establishes a smooth path between the two quantum many-body states. In Bachmann et al. (2012), it is show that such a path can be realized with a smoothly parametrized unitary $U(s)$ whose generator $G(s)$ is also a local operator, $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$, and naturally $\rho_{1}=U(s=1)\rho_{0}U^{\dagger}(s=1)$. This is further elaborated in Appendix B.
@@
With the Nielsen’s QCC of geometrically local quantum circuit as a distance measure, the proposed theoretically optimal kernel for the unsupervised learning of topological order is given by
@@
up to a normalization constant, where $\beta>0$ is a hyperparameter. The unitarity of the quantum circuit indicates that Nielsen’s QCC is symmetric under permutations of the two quantum states, i.e., $\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{N}}\left(\rho_{1}\rightarrow\rho_{0}\right)$, which leads to a symmetric kernel in (3).
III Unsupervised machine learning of topological order derived from quantum circuit complexity

### v2

@@ Abstract
Inspired by the close relationship between Kolmogorov complexity and unsupervised machine learning, we explore quantum circuit complexity, an important concept in quantum computation and quantum information science, as a pivot to understand and to build interpretable and efficient unsupervised machine learning for topological order in quantum many-body systems. We argue that Nielsen’s quantum circuit complexity represents an intrinsic topological distance or similarity measure between topological quantum many-body phases of matter, and as such plays a central role in interpretable manifold learning of topological order. To span a bridge from conceptual power to practical applicability, we present two theorems that connect Nielsen’s quantum circuit complexity for the quantum path planning between two arbitrary quantum many-body states with quantum Fisher complexity (Bures distance) and entanglement generation, respectively. Leveraging these connections, fidelity-based and entanglement-based similarity measures or kernels, which are more practical for implementation, are formulated. Using the two proposed distance measures, unsupervised manifold learning of quantum phases of the bond-alternating XXZ spin chain, the ground state of Kitaev’s toric code and random product states, is conducted, demonstrating their superior performance. Moreover, we find that the entanglement-based approach, which captures the long-range structure of quantum entanglement of topological orders, is more robust to local Haar random noises. Relations with classical shadow tomography and shadow kernel learning are also discussed, where the latter can be naturally understood from our approach. Our results establish connections between key concepts and tools of quantum circuit computation, quantum complexity, quantum metrology, and machine learning of topological quantum order.
@@ I Introduction
Recent progress has shown the potential of machine learning for classifying topological phases of matter Carleo and Troyer (2017); Carrasquilla and Melko (2017); Van Nieuwenburg et al. (2017); Zhang and Kim (2017); Liu and van Nieuwenburg (2018); Zhang et al. (2018a); Sun et al. (2018); Carleo et al. (2019); Ming et al. (2019); Rem et al. (2019); Lian et al. (2019); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020); Long et al. (2020); Greplova et al. (2020); Balabanov and Granath (2020); Yu and Deng (2021); Mendes-Santos et al. (2021); Yu et al. (2022); Leykam and Angelakis (2023); Smolina et al. (2024); Yu et al. (2024); Das Sarma et al. (2019). In particular, the supervised approach to quantum many-body physics problems, which can resort to well established theories of learnability and generalizability Valiant (1984); Mohri et al. (2018); Jacot et al. (2018), has resulted in the formulation of rigorous guarantees Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). However, in many practical scenarios, where a priori knowledge of the different phases of matter (and accordingly their labels) does not exist, only unsupervised machine learning can provide access to the desired information. Despite the practical relevance of unsupervised learning, a rigorous theory, tailored to topological quantum order and offering both interpretability and generalizability, has not yet been fully explored.
@@
One prevalent approach to unsupervised learning is manifold learning Tenenbaum et al. (2000); Roweis and Saul (2000); Maaten and Hinton (2008), where the data is assumed to be distributed on a manifold characterized by an intrinsic distance metric. It is followed by a nonlinear mapping of the data manifold into a low-dimensional Euclidean space while preserving the original similarities and structures. Among many manifold-learning algorithms, diffusion map Coifman et al. (2005); Nadler et al. (2006), t-SNE Maaten and Hinton (2008), and kernel principal component analysis (PCA) Schölkopf et al. (1998); Mika et al. (1999) require a kernel function constructed from the intrinsic distance. The kernel can be used either to build a probabilistic graph representation of the data manifold (e.g., in the diffusion map or the t-SNE), or for PCA in the nonlinear feature space (e.g., in the kernel PCA). Alternatively, the intrinsic distance metric itself can be applied directly to non-kernel manifold learning, such as the Isomap Tenenbaum et al. (2000) and the metric-multidimensional scaling (metric-MDS) Demaine et al. (2021); Skl (a). In the context of topological quantum order, the basic task then is to formulate a distance metric that encodes the nonlinear and non-Euclidean features of topological quantum states and represents topological similarities or distances by exploiting the primitive notion of topological equivalence. Topologically equivalent quantum states, by definition, can be continuously transformed into each other via a smooth unitary transformation generated by a local Hermitian operator with bounded operator norm Zeng et al. (2015). In this spirit, a similarity measure based on a path-finding algorithm has been proposed in Scheurer and Slager (2020) and has shown good performance in clustering symmetry-protected band topologies, which are characterized by short-range quantum entanglement. Other topological similarity measures, that focus on the closing of the spectral gap at topological transitions, have also been proposed and demonstrated Che et al. (2020); Long and Zhang (2023).
@@
The inherent problems of previous manifold-learning approaches call for a new similarity ansatz, an ansatz which is able to capture the topological distance and entanglement patterns at different scales while at the same time remaining interpretable and efficiently computable (see Fig. 1 for a schematic of the current state regarding the application of machine learning to the classification of topological phases of matter). To this end, we here adopt the alternative but equivalent perspective that two topologically equivalent multi-qubit states can be transformed into each other via a constant-depth quantum circuit composed of geometrically local quantum gates Zeng et al. (2015); Zeng and Wen (2015), i.e., the smooth unitary transformation that connects them can be decomposed into a shallow quantum circuit. This implies that it is algorithmically trivial or cheap to generate one quantum state given the access to another in the same topological class by fully exploiting the shared structure and entanglement patterns between them.
@@
Figure 1: Landscape of machine learning topological phases of matter. Supervised learning requires prior information or labels, whose efficiency guarantees have been extensively investigated Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). Under the constraint of geometric locality Lewis et al. (2024); Rouzé et al. (2024) or with a bounded number (No.) of parameters (param.) Che et al. (2024), provably efficient machine learning of quantum many-body systems has been established, e.g., in the framework of probably approximately correct (PAC)-learnable Valiant (1984); Che et al. (2024). Supervised learning is conceptually more straightforward due to the existence of well-established statistical learning theories. In contrast, unsupervised learning, whose theory is more complicated and less addressed, becomes practically relevant when there is no access to prior labels. The symmetry-protected topological (SPT) order features short-range entanglement, while the topological quantum order exhibits long-range entanglement. Unsupervised learning with the path-finding algorithm Scheurer and Slager (2020) and kernels focusing on closing gaps at topological transitions Che et al. (2020); Long and Zhang (2023), respectively, have shown good performance and interpretability for SPT orders or band topologies. However, the direct generalization of these methods to more complex quantum systems with increased complexity and quantum entanglement faces challenges, due to the intrinsic hardness of identifying (or even the undecidability Cubitt et al. (2015) of) the spectral gap for generic quantum many-body Hamiltonians (See more elaborations and references in the introduction of the main text). Note that reinforcement learning for quantum systems are not included in this landscape.
@@
The correspondence between topological equivalence of quantum states and shallow quantum circuit transformations indicates that, a complexity measure may be related to the topological distance. This is similar to the situation with Kolmogorov complexity, which represents an informational distance between two strings or objects Bennett et al. (2010), and may offer a basis for theoretically optimal solution to unsupervised machine learning Ily (See more details in the following section). We then assert that, the (conditional) generation and prediction of quantum states with the minimal quantum circuit cost may offer a theoretically optimal solution to the unsupervised learning of topological order. The complexity of the generating circuit can be interpreted as a quantum informational distance measure that captures the entanglement change at different scales. Specifically, we propose to use Nielsen’s quantum circuit complexity (QCC) of the unitary path connecting two quantum states as a suitable distance measure that can be used for the unsupervised manifold learning of topological order. While the QCC is already a central concept in quantum information theory Chi-Chih Yao (1993); Brandão et al. (2021); Haferkamp et al. (2022); Yi et al. (2024); Liu et al. (2020a), quantum field theory Chapman et al. (2018); Hackl and Myers (2018); Khan et al. (2018) and quantum gravity Susskind (2014); Stanford and Susskind (2014); Brown et al. (2016a, b, 2023), its connection with, and relevance for, the unsupervised machine learning of topological phases of matter has remained mostly unexplored.
@@
In this work, we use the QCC as a pivotal tool to understand and construct unsupervised machine learning of topological order for quantum many-body systems, in an interpretable and efficient manner. Although the exact QCC is generically intractable, we identify various quantities which are upper bounded by the QCC while preserving useful information and remaining implementable, and which thus can be efficiently utilized for classifying topological quantum many-body phases. Specifically, we establish entanglement-based and fidelity-based similarities / kernels, respectively. Numerical experiments with a bond-alternating XXZ qubit chain and Kitaev’s toric code model, respectively, demonstrate the excellent performance of these methods.
II Results
II.1 Quantum path planning, topological order and quantum circuit complexity
@@
As mentioned above, unsupervised machine learning of band topology has been extensively investigated. Band topologies, or the more general symmetry-protected topological (SPT) orders, such as those exhibited by topological insulators, are characterized by the absence of long-range quantum entanglement. This implies that, if the constraint of preserving the symmetry is dropped, SPT quantum states can be transformed smoothly into product states. In contrast, topological quantum order refers to quantum states with long-range quantum entanglement Zeng et al. (2015); for instance, the ground state of Kitaev’s toric code exhibits topological quantum order that manifests in non-zero topological entanglement entropy Kitaev and Preskill (2006); Levin and Wen (2006) and global logical operations Kitaev (2003), while SPT phases are confined to short-range entanglement. Thus, the task of identifying topological quantum order with unsupervised learning presupposes to construct similarity measures with feature maps that capture quantum entanglement at a variety of scales. For this purpose, we start by defining the quantum path planning between two multi-qubit quantum states, which lays the ground for our search for a distance measure of clustering topological orders.
@@
[Quantum path planning (QPP).] The QPP problem is defined as: Given two arbitrary pure $n$-qubit quantum states $\rho_{0}$ and $\rho_{1}$, find continuous paths $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$ $(s\in[0,1])$ that connect $\rho_{0}$ and $\rho_{1}$, where the generator $G(s)$ of $U(s)$ is a hermitian operator, with the operator norm $\|\partial_{s}U(s)\|_{\infty}=\|G(s)\|_{\infty}<+\infty$ $(\forall s\in[0,1])$. In general, we have that $U(s)\in\mathrm{SU}\left(2^{n}\right)$ and $iG(s)\in\mathfrak{su}\left(2^{n}\right)$, with the unitary given by
@@
where $U(s=0)=\mathrm{Id}$ is the identity and $\cal{P}$} denotes the path ordering operator. In practical implementation, the QPP can involve finding the optimal path under constraints.
@@
(Topological equivalence Zeng and Wen (2015); Zeng et al. (2015).) We say that two gapped ground states $\rho_{0}$ and $\rho_{1}$ of local Hamiltonians $H_{0}$ and $H_{1}$, respectively, are in the same topological phase if there is a local unitary that maps the two states into each other without closing the energy gap. In other words, there is a smooth path of local Hamiltonians $H(s)$ that connects the two parent Hamiltonians with $H(s=0)=H_{0}$ and $H(s=1)=H_{1}$ for $s\in[0,1]$. The Hamiltonian $H(s)$ is gapped above its ground state $\rho(s)$ for all $s\in[0,1]$, where $\rho(s)$ smoothly connects $\rho_{0}$ and $\rho_{1}$, which establishes a smooth path between the two quantum many-body states. In Bachmann et al. (2012), it is show that such a path can be realized with a smoothly parametrized unitary $U(s)$ whose generator $G(s)$ is also a local operator, $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$, and naturally $\rho_{1}=U(s=1)\rho_{0}U^{\dagger}(s=1)$. This is further elaborated in Sec. II of the Supplementary Information.
@@
With the Nielsen’s QCC of geometrically local quantum circuit as a topological and informational distance measure, the proposed theoretically optimal similarity/ kernel function for the unsupervised manifold learning of topological order is given by
@@
up to a normalization constant, where $\beta>0$ is a hyperparameter. The unitarity of the quantum circuit indicates that Nielsen’s QCC is symmetric under permutations of the two quantum states, i.e., $\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{N}}\left(\rho_{1}\rightarrow\rho_{0}\right)$, which leads to a symmetric kernel in (3). The similarity function in (3) can be used to construct conditional probabilities in the diffusion map Coifman et al. (2005); Nadler et al. (2006) and the t-SNE Maaten and Hinton (2008) manifold learning, or can be used for kernel PCA Schölkopf et al. (1998); Mika et al. (1999) in the kernel feature space. Physically, Nielsen’s QCC can be viewed as a global order parameter of topological quantum orders, and also serves as an ideal topological distance in non-kernel manifold learning such as the metric-MDS Demaine et al. (2021); Skl (a).
II.2 Unsupervised machine learning of topological order derived from quantum circuit complexity